# One Patch to Caption Them All: A Unified Zero-Shot Captioning Framework

### v2

@@ 3.2 From Patches to Regions
Dense Captioning requires locating salient regions in an image and generating their descriptions. Following defined evaluation protocols (Johnson et al., 2016), we focus on captioning already defined boxes, effectively removing the localization subtask, which can be tackled using additional region-proposal models. Given a bounding box $B$ and the set $S_{B}$ of indexes of patches that intersect with $B$, we obtain the representation of the region $\mathbf{v}_{B}=\operatorname*{avg}_{i\in S_{B}}(\mathbf{v}_{i})$.
@@ 4.1 Datasets and Metrics
We build a benchmark for Trace Captioning exploiting Localized Narratives (Pont-Tuset et al., 2020b) — a dataset in which annotators vocally described objects in images while moving the mouse pointer over the described object. The dataset provides temporal annotated voice transcriptions and mouse traces for the images of many standard captioning datasets. We took the labeled COCO (Lin et al., 2014; Chen et al., 2015) with splits defined in Karpathy & Fei-Fei (2015) and Flickr30K [^1] Licenses for COCO and Visual Genome are CC-BY 4.0. Flickr30k images are under the terms of Flickr’s user agreement.(Young et al., 2014) test subsets to build the trace captioning evaluation datasets. We split long traces and transcriptions for each image into sentences, and we discard the parts of the traces that are not temporally located between the start and the end of each sentence. We discarded noisy sentences — such as the ones describing image properties (The image is blurred, the image is edited, …) and rewrote each sentence removing uncertainties typical of voice descriptions in a more concise and caption-like style through a few-shots prompted LLM (LLama 3 Dubey et al. (2024)). After annotation cleaning, 51 COCO images resulted without clean sentences and were discarded. Each sub-trace and relative sentence comprise an independent sample, that is, we ignore the temporal information of consecutive sentences and focus on evaluating the description of each sub-trace only. Samples and additional details are available in §10.
@@
We assess the performance on dense captioning tasks following the evaluation procedure of Johnson et al. (2016), omitting the bounding box proposal and evaluating only the bounding box captioning task, using ground-truth boxes as input for the models. In addition to standard caption metrics, for this task, we also report the mAP as originally defined by Johnson et al. (2016). We use the Visual Genome (VG) v1.2  (Johnson et al., 2016; Krishna et al., 2017) and VG-COCO test splits (Li et al., 2019). The former comprises 5000 images from VG, while the latter contains 2476 images present in both VG and COCO. Both contain multiple bounding box annotations per image with descriptions.
@@
We follow the evaluation protocol of Cornia et al. (2019) that originally introduced region-set captioning. We use the Flickr30K Entities (Plummer et al., 2015) and the COCO Entities (Cornia et al., 2019) datasets. Each record comprises an image, a set of bounding boxes of variable length, and a ground-truth controlled caption. We evaluate on the test splits, comprising of images in the Karpathy & Fei-Fei (2015) test splits, that consist of 3569 and 1000 images for COCO and Flickr30k versions, respectively.
@@
We follow the standard evaluation pipeline for zero-shot image captioning, generating captions for the 5000 images in Karpathy’s COCO test split. We compare with these state-of-the-art models: DeCap (Li et al., 2023b), CLOSE (Gu et al., 2023), ZeroCap (Tewel et al., 2022), MAGIC (Su et al., 2022), ViECap (Fei et al., 2023), CapDec (Nukrai et al., 2022), EntroCap (Yan et al., 2025), MeaCap [^2] We picked the MeaCapInvLM, which uses a GPT-2 decoder and achieves the highest scores in the zero-shot captioning task. (Zeng et al., 2024), and MERCap (Zeng et al., 2025). For DeCap, ViECap, and MeaCap, we also report the best results reproduced by us.
@@
All datasets used in our evaluation provide a ground-truth caption for each annotation. We therefore adopt standard captioning metrics to assess the similarity between generated and reference captions. In particular, we focus on CIDEr (C) (Vedantam et al., 2015), which captures syntactic overlap, and RefPAC Score (P) (Sarto et al., 2023), a more recent metric that quantifies semantic similarity independently of caption phrasing. For completeness, results with other traditionally used metrics — BLEU@4 (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), ROUGE-L (Lin, 2004), and SPICE (Anderson et al., 2016) — are reported in supplementary materials, as they follow the same trends. For the image captioning task, we additionally report CLIP-Score (Hessel et al., 2021), which measures the alignment between an image and its generated caption in the joint CLIP vision–language space. This metric is not applicable to region-set, trace, or dense captioning tasks, where captions describe local regions rather than the entire image.
@@ 4.2 Backbone Selection
The choice of the visual backbone is crucial for our patch-centric framework, as the quality and semantic richness of the patch features directly impact the captioning performance. We tested several state-of-the-art vision-language models pre-trained without region-level supervision and evaluated their effectiveness within the framework. We tested vanilla CLIP (Radford et al., 2021), three CLIP adaptations for dense tasks — DenseCLIP (Rao et al., 2022), INViTE (Chen et al., 2024), and ProxyCLIP (Lan et al., 2024) —, and two methods with visual encoders based on DINOv2 (Oquab et al., 2023)DINO.txt (Jose et al., 2025) and Talk2DINO (Barsellotti et al., 2024).
Patches are aggregated as described in §3.2, while for the zero-shot decoder, we align with the setting of Li et al. (2023b), and use a prefix GPT-2 style decoder (SM§6 reports implementation details), with a memory-based latent projection approach as mitigation strategy for handling the modality gap. Specifically, before decoding, the region representation $\mathbf{v}$ is projected into the text embedding space as a similarity-weighted linear combination of memory elements, $\mathbf{v}_{\text{proj}}=M\,\alpha$ with $\alpha=\mathrm{softmax}\!\left(\tfrac{1}{\tau}M^{\top}\mathbf{v}\right)$, where $M=[\mathbf{m}_{1},\dots,\mathbf{m}_{N}]$ stores the text embeddings $\mathbf{m}_{j}=\psi_{t}(t_{j})$ and $\tau>0$ controls the sharpness of the weighting distribution. This choice enables us to be directly comparable with other works using the same architecture in the all the following experiments, besides also performing the best among the tested zero-shot decoder methods (see SM§9). In SM§8, we also study how the choice of the memory bank $M$ affects the captioning performance, providing an upper bound to metrics.
Captioning Task:
Trace
Dense
(VG v1.2)
Region-Set
(COCO Entities)
Image
Encoder Backbone C P C P C P C P CLIP-S
CLIP CLIP B/16 10.9 75.0 10.9 74.2 41.6 78.8 42.1 84.0 66.2
DenseCLIP CLIP B/16 18.6 75.3 19.9 75.2 51.0 77.6 28.0 77.0 57.3
INViTE CLIP B/16 13.8 76.4 16.8 77.3 43.3 78.9 21.3 79.1 60.6
ProxyCLIP DINO B/8 + CLIP B/16 16.7 75.7 15.7 76.0 41.2 78.4 28.7 79.0 61.7
ProxyCLIP DINOv2 B/14 + CLIP B/16 16.5 75.7 15.5 76.0 40.6 78.5 27.4 78.6 61.0
DINO.txt DINOv2 B/14 23.2 78.8 23.4 78.5 91.8 86.3 67.8 87.2 70.8
Talk2DINO DINOv2 B/14 27.9 78.7 31.9 78.8 109.1 87.5 69.2 87.4 72.8
@@
Table 1 shows that backbone effectiveness in our framework is closely tied to capturing fine-grained local semantics. Standard CLIP performs poorly, indicating that its patch tokens lack the spatial detail needed for our tasks (Mukhoti et al., 2023; Ranasinghe et al., 2023; Bica et al., 2024). Backbones that strengthen CLIP local representations, such as INViTE and DenseCLIP, achieve stronger results, supporting our hypothesis. The best performance comes from DINOv2-based models, including DINO.txt and Talk2DINO, with the latter emerging as the most effective encoder. This underscores the importance of semantically rich patch-level features for high-quality region-level captions. For this reason, we show results using Talk2DINO as the default backbone in subsequent experiments.
@@ 4.3 Comparison with SOTA
Despite significant advances in zero-shot image captioning, we are unaware of any prior methods specifically tailored for zero-shot regional captioning tasks. Existing zero-shot captioners are usually evaluated only at the level of whole images, without any mechanisms to natively attend to arbitrary regions. To rigorously quantify the benefit of our Patch-ioner framework, in Table 2, we compare against both state-of-the-art zero-shot image captioners and adapted baselines: (i) state-of-the-art whole-image zero-shot captioners in their standard setting or applied to region crops, simulating regional captioning by isolating local content, and (ii) region-supervised encoders, thus outside our no-region-label setting, that leverage mask-based (AlphaCLIP, Sun et al. (2024)) or crop-based (RegionCLIP, Zhong et al. (2022)) attention coupled with the same zero-shot decoder, allowing them to attend to specific regions. This design ensures that our evaluation covers the strongest available baselines for both image-level and region-level zero-shot captioning. Note that we do not compare with large multimodal models tackling regional understanding and captioning, such as Guo et al. (2024); Hua et al. (2025), as they are outside of our assumptions on the available supervision by training on region-level data.
In addition to our strongest model (i.e., Talk2DINO with the memory-based mitigation strategy), we also report other combinations that express existing zero-shot captioning approaches (CLOSE (Gu et al., 2023), CapDec (Nukrai et al., 2022), VieCap (Fei et al., 2023), and MeaCap (Zeng et al., 2024)) but replacing the original CLIP backbone. Specifically, CLOSE and CapDec can be expressed by choosing the noise-injection mitigation strategy and using the standard decoder pipeline described in §4.2. The same applies to ViECap and MeaCap, although with the addition of extra knowledge in the text decoding step: ViECap uses external entity-aware prompts, while MeaCap leverages also structured concept retrieval from a knowledge base.
We use two datasets for each task — a COCO-derived dataset and an additional dataset such as Visual Genome (Krishna et al., 2017) or Flickr30k (Young et al., 2014). Figure 6 shows qualitative results.
@@
In the trace and dense captioning tasks, which emphasize local visual content, our patch-centric framework significantly outperforms all baselines across metrics. For trace captioning (Table 2, first group), our patch-based formulation outperforms whole-image captioners and crop-based adaptations. Models relying on global CLS representations fail to capture the precise objects and attributes under the trace. Even AlphaCLIP, that is a region-supervised backbone that can be naturally applied to traces, lags behind, underlying intrinsic limitations of the pretrained CLIP backbone. Dense captioning shows a similar trend (Table 2, second group), with our models outperforming baselines. For this task, we report crop-based adaptations for DeCap, ViECap, and MeaCap, as they provide a stronger baseline with respect to the same models applied to the global CLS (see Table 4 in SM). Although isolating regional content, these models discard broader contextual cues that are crucial for coherent dense descriptions.
@@
On region-set captioning (Table 2, third group), our models once again achieves state-of-the-art results, outperforming both zero-shot baselines and region-supervised models. Note that the region-set captioning task tends to align more closely with image-level captioning rather than strictly focusing on localized regions (see Figure 6 and Figure 7 in SM), as regions are intended to control an image-level caption [^3] This is expected since the ground-truth captions in the COCO Entities dataset originate from the image-level annotations of COCO, as stated by Cornia et al. (2019).. Thus, global models also tend to perform well on those tasks, showing a narrower gap with respect to our new state of the art. By aggregating patch embeddings from arbitrary and possibly disjoint sets of regions, the model produces coherent and contextually rich captions that align with the collective semantics of the chosen areas. In contrast, whole-image methods, including ViECap and DeCap, cannot naturally incorporate regional cues, limiting their effectiveness in this setting. Importantly, our patch-based aggregation even surpasses AlphaCLIP, which relies on explicit mask supervision.
@@
In whole-image captioning (Table 2, fourth group), our results remain competitive with the strongest zero-shot captioners but are slightly behind dedicated image-centric architectures such as MERCap (Zeng et al., 2025) and EntroCap (Yan et al., 2025). For this task, we report the performance of our models using an attention-based weighting, when helpful, as it usually performs marginally better than the standard average patch aggregation and thus represents the best available model for this task and reaches the smaller gap with state-of-the-art models (see SM§8). Notably, adding structured external knowledge or filtered retrieval (as in ViECap or MeaCap) on noise-based decoders improves fluency and informativeness, suggesting such modules are complementary than substitutes for strong regional semantics. However, they compare similarly to the model using the memory-based decoder.
@@
ZeroCap (Tewel et al., 2022) CVPR’22 - - - - - - - - - - - - - - 14.6 - - - - -
MAGIC (Su et al., 2022) ArXiv’22 - - - - - - - - - - - - - - 49.3 - - 17.5 - -
CLOSE (Gu et al., 2023) ICCV’23 - - - - - - - - - - - - - - 81.2 - - - - -
CapDec (Nukrai et al., 2022) EMNLP’22 - - - - - - - - - - - - - - 91.8 - - 35.7 - -
EntroCap (Yan et al., 2025) NEUCOM’25 - - - - - - - - - - - - - - 94.3 - - 41.5 - -
MERCap (Zeng et al., 2025) AAAI’25 - - - - - - - - - - - - - - 96.0 - - 45.6 - -
ViECap (Fei et al., 2023) ICCV’23 24.3 74.4 12.0 68.8 14.90 26.4 74.3 15.2 26.6 74.3 102.7 85.0 31.8 74.9 89.7 88.5 75.6 29.8 80.6 70.5
MeaCap (Zeng et al., 2024) CVPR’24 22.5 74.4 12.6 69.8 15.01 28.6 75.1 16.0 28.9 75.1 97.9 85.2 38.6 76.4 86.0 88.6 77.8 40.1 82.8 73.9
DeCap (Li et al., 2023b) ICLR’23 20.5 75.3 11.2 71.0 17.75 24.6 77.8 17.8 24.9 77.7 95.1 87.4 39.4 78.8 87.4 90.6 79.3 40.0 84.8 76.3
@@
RegionCLIP (Zhong et al., 2022) CVPR’22 + Mem. ($\simeq$ DeCap) - - - - 15.85 21.7 76.7 16.01 21.0 75.4 - - - - 93.4 91.2 77.5 38.8 84.5 73.6
AlphaCLIP (Sun et al., 2024) CVPR’24 + Mem. ($\simeq$ DeCap) 21.3 75.4 11.8 71.0 14.63 19.1 73.9 14.82 19.4 73.8 95.1 87.4 39.5 78.8 89.7 91.1 78.2 40.9 85.4 75.0
@@
Table 2: Comparison of our Patch-ioner framework, using Talk2DINO (T2D), with state-of-the-art zero-shot captioning methods on trace, dense, region-set, and image captioning tasks. Our approach consistently outperforms whole-image and region-supervised baselines in local, fine-grained captioning tasks, while achieving competitive results on whole-image captioning. The table reports CIDEr (C), RefPAC (P), mean average precision (mAP) for dense captioning, and CLIP-Score (CLIP-S) when applicable; best and second-best results are highlighted in bold and underline, respectively.
@@
Figure 2: Qualitative results from finer (left) to coarser (right) tasks. Note the discrepancy of predicted and ground-truth captions when an image-level (DeCap, DeCap (Crop)) or a CLIP-based regional (CLIP + Mem.) captioner is applied, with respect to our Talk2DINO-based model.

### v3

@@ 3.2 From Patches to Regions
Dense Captioning requires locating salient regions in an image and generating their descriptions. Following defined evaluation protocols [23], we focus on captioning already defined boxes, effectively removing the localization subtask, which can be tackled using additional region-proposal models. Given a bounding box $B$ and the set $S_{B}$ of indexes of patches that intersect with $B$, we obtain the representation of the region $\mathbf{v}_{B}=\operatorname*{avg}_{i\in S_{B}}(\mathbf{v}_{i})$.
@@ 4.1 Datasets and Metrics
We build a benchmark for Trace Captioning exploiting Localized Narratives [49] — a dataset in which annotators vocally described objects in images while moving the mouse pointer over the described object. The dataset provides temporal annotated voice transcriptions and mouse traces for the images of many standard captioning datasets. We took the labeled COCO [38, 9] with splits defined in [26] and Flickr30K[66] test subsets to build the trace captioning evaluation datasets. We split long traces and transcriptions for each image into sentences, and we discard the parts of the traces that are not temporally located between the start and the end of each sentence. We discarded noisy sentences — such as the ones describing image properties (The image is blurred, the image is edited, …) and rewrote each sentence removing uncertainties typical of voice descriptions in a more concise and caption-like style through a few-shots prompted LLM (LLama 3 [13]). After annotation cleaning, 51 COCO images resulted without clean sentences and were discarded. Each sub-trace and relative sentence comprise an independent sample, that is, we ignore the temporal information of consecutive sentences and focus on evaluating the description of each sub-trace only. The final benchmark includes 4,949 COCO images with 14,283 captions (average caption length: 7.8 words) and 1,000 Flickr30K images with 26,614 captions (average caption length: 6.7 words). Samples and additional details are available in §13.
@@
We assess the performance on dense captioning tasks following the evaluation procedure of [23], omitting the bounding box proposal and evaluating only the bounding box captioning task, using ground-truth boxes as input for the models. In addition to standard caption metrics, for this task, we also report the mAP as originally defined by [23]. We use the Visual Genome (VG) v1.2  [23, 28] and VG-COCO test splits [34]. The former comprises 5000 images from VG, while the latter contains 2476 images present in both VG and COCO. Both contain multiple bounding box annotations per image with descriptions.
@@
We follow the evaluation protocol of [10] that originally introduced region-set captioning. We use the Flickr30K Entities [47] and the COCO Entities [10] datasets. Each record comprises an image, a set of bounding boxes of variable length, and a ground-truth controlled caption. We evaluate on the test splits, comprising images from the Karpathy [26] splits, which consist of 3569 and 1000 images for COCO and Flickr30k versions, respectively.
@@
We follow the standard evaluation for zero-shot captioners, generating captions for the 5000 images in Karpathy’s COCO test split. We compare with DeCap [33], CLOSE [18], ZeroCap [59], MAGIC [56], ViECap [15], CapDec [43], EntroCap [63], MeaCap [^1] We picked the MeaCapInvLM, which uses a GPT-2 decoder and achieves the highest scores in the zero-shot captioning task. [68], MERCap [67], IFCap [31], and Diffusion Bridge [30].
@@
All datasets used in our evaluation provide a ground-truth caption for each annotation. We therefore adopt standard captioning metrics to assess the similarity between generated and reference captions. In particular, we focus on CIDEr (C) [61], which captures syntactic overlap, and RefPAC Score (P) [54], a more recent metric that quantifies semantic similarity independently of caption phrasing. For completeness, results with other traditionally used metrics — BLEU@4 [46], METEOR [4], ROUGE-L [37], and SPICE [2] — are reported in supplementary materials, as they follow the same trends. For the image captioning task, we additionally report CLIP-Score [20], which measures the alignment between an image and its generated caption in the joint CLIP space. This metric is not applicable to region-set, trace, or dense captioning tasks, where captions describe local regions rather than the entire image.
@@ 4.2 Backbone Selection
The choice of the visual backbone is crucial for our patch-centric framework, as the quality and semantic richness of the patch features directly impact the captioning performance. We tested several state-of-the-art vision-language models pre-trained without region-level supervision and evaluated their effectiveness within the framework. We tested vanilla CLIP [50], three CLIP adaptations for dense tasks — DenseCLIP [53], INViTE [8], and ProxyCLIP [29] —, SigLIP2 [60], and two methods with visual encoders based on DINOv2 [44]DINO.txt [25] and Talk2DINO [5]. In §7 we provide further details on each backbone adopted.
Patches are aggregated as described in §3.2, while for the zero-shot decoder, we align with the setting of [33], and use a prefix GPT-2 style decoder (SM§6 reports implementation details), with a memory-based latent projection approach as mitigation strategy for handling the modality gap. Specifically, before decoding, the region representation $\mathbf{v}$ is projected into the text embedding space as a similarity-weighted linear combination of memory elements, $\mathbf{v}_{\text{proj}}=M\,\alpha$ with $\alpha=\mathrm{softmax}\!\left(\tfrac{1}{\tau}M^{\top}\mathbf{v}\right)$, where $M=[\mathbf{m}_{1},\dots,\mathbf{m}_{N}]$ stores the text embeddings $\mathbf{m}_{j}=\psi_{t}(t_{j})$ and $\tau>0$ controls the sharpness of the weighting distribution. This choice enables us to be directly comparable with other works using the same architecture in the all the following experiments, besides also performing the best among the tested zero-shot decoder methods (see SM§12). In SM§11, we also study how the choice of the memory bank $M$ affects the captioning performance, providing an upper bound to metrics.
Captioning Task: (Dataset) Trace (COCO) Dense (VG v1.2) Region-Set (COCO Entities) Image (COCO)
Vision Backbone Text Backbone C P C P C P C P CLIP-S
CLIP CLIP B/16 CLIP B/16 10.9 75.0 10.9 74.2 41.6 78.8 42.1 84.0 66.2
DenseCLIP CLIP B/16 CLIP B/16 18.6 75.3 19.9 75.2 51.0 77.6 28.0 77.0 57.3
INViTE CLIP B/16 CLIP B/16 13.8 76.4 16.8 77.3 43.3 78.9 21.3 79.1 60.6
ProxyCLIP DINO B/8 + CLIP B/16 CLIP B/16 16.7 75.7 15.7 76.0 41.2 78.4 28.7 79.0 61.7
ProxyCLIP DINOv2 B/14 + CLIP B/16 CLIP B/16 16.5 75.7 15.5 76.0 40.6 78.5 27.4 78.6 61.0
SigLIP2 SigLIP2 B/16 SigLIP2 B/16 18.3 73.6 19.8 73.4 47.2 76.7 27.7 77.0 56.4
DINO.txt DINOv2 B/14 DINO.txt 23.2 78.8 23.4 78.5 91.8 86.3 67.8 87.2 70.8
Talk2DINO DINOv2 B/14 Talk2DINO 27.9 78.7 31.9 78.8 109.1 87.5 69.2 87.4 72.8
@@
Table 1 shows that backbone effectiveness in our framework is closely tied to capturing fine-grained local semantics. Standard CLIP performs poorly, indicating that its patch tokens lack the spatial detail needed for our tasks [41, 52, 6]. Backbones that strengthen CLIP local representations, such as INViTE and DenseCLIP, achieve stronger results, supporting our hypothesis. The best performance comes from DINOv2-based models, including DINO.txt and Talk2DINO, with the latter emerging as the most effective encoder. This underscores the importance of semantically rich patch-level features for high-quality region-level captions. For this reason, we show results using Talk2DINO as the default backbone in subsequent experiments.
@@ 4.3 Comparison with SOTA
Despite significant advances in zero-shot image captioning, we are unaware of any prior methods specifically tailored for zero-shot regional captioning tasks. Existing zero-shot captioners are usually evaluated only at the level of whole images, without any mechanisms to natively attend to arbitrary regions. To rigorously quantify the benefit of our approach, in Table 2, we compare against both state-of-the-art zero-shot image captioners and adapted baselines: (i) whole-image zero-shot captioners in their standard setting or applied to region crops, simulating regional captioning by isolating local content, and (ii) region-supervised encoders, thus outside our no-region-label setting, that leverage mask-based (AlphaCLIP, [57]) or crop-based (RegionCLIP, [71]) attention coupled with the same zero-shot decoder, allowing them to attend to specific regions. This design ensures that our evaluation covers the strongest available baselines for both image-level and region-level zero-shot captioning. While we exclude models trained on region-level data [19, 21], in §8 we compare our framework against representative LMMs. Despite their massive end-to-end pre-training on image-text pairs, our text-only method achieves superior performance on fine-grained regional tasks.
In addition to our strongest model (i.e., Talk2DINO with the memory-based mitigation strategy), we also report other combinations that express existing zero-shot captioning approaches (CLOSE [18], CapDec [43], VieCap [15], MeaCap [68] and Diffusion Bridge [30]) but replacing the original CLIP backbone. Specifically, CLOSE and CapDec can be expressed by choosing the noise-injection mitigation strategy and using the standard decoder pipeline described in §4.2. The same applies to ViECap and MeaCap, although with the addition of extra knowledge in the text decoding step: ViECap uses external entity-aware prompts, while MeaCap leverages also structured concept retrieval from a knowledge base. Finally, Diffusion Bridge is implemented by applying a diffusion model trained on the textual manifold to the patch-level representation.
We use two datasets for each task — a COCO-derived dataset and an additional dataset such as Visual Genome [28] or Flickr30k [66]. Figure 7 shows qualitative results.
@@
In trace and dense captioning tasks, which emphasize local visual content, our patch-centric framework significantly outperforms all baselines across metrics. In trace captioning (Table 2, 1st group), our formulation outperforms image-based captioners and crop-based adaptations. Models relying on global CLS representations fail to capture the precise objects and attributes under the trace. Even AlphaCLIP, a region-supervised backbone applicable to traces, lags behind, underlying intrinsic limitations of the pretrained CLIP backbone. Dense captioning shows a similar trend (Table 2, 2nd group), with our models outperforming baselines. For this task, we report crop-based adaptations for DeCap, ViECap, MeaCap, IFCap, and Diffusion Bridge, as they provide a stronger baseline with respect to the same models applied to the global CLS (see Table 7 in SM). Although isolating regional content, these models discard broader contextual cues that are crucial for coherent dense descriptions.
@@
Also in region-set captioning (Table 2, 3rd group), our models achieves state-of-the-art results, outperforming both zero-shot baselines and region-supervised models. Region-set captioning tends to align more closely with image-level captioning rather than strictly focusing on localized regions (see Figure 7 and Figure 8 in SM), as regions are intended to control an image-level caption [^2] This is expected since the ground-truth captions in the COCO Entities dataset originate from the image-level annotations of COCO [10].. Thus, global models also tend to perform well on those tasks, showing a narrower gap with respect to ours. By aggregating patch embeddings from arbitrary and possibly disjoint sets of regions, the model produces coherent and contextually rich captions that align with the collective semantics of the chosen areas. In contrast, whole-image methods cannot naturally incorporate regional cues, limiting their effectiveness in this setting. Importantly, our patch-based aggregation even surpasses AlphaCLIP, which relies on explicit mask supervision.
@@
In whole-image captioning (Table 2, 4th group), our results remain competitive with the strongest zero-shot captioners but are slightly behind dedicated image-centric architectures such as MERCap [67] and EntroCap [63]. For this task, we report the performance of our models using an attention-based weighting, when helpful, as it usually performs marginally better than the standard average patch aggregation and thus represents the best available model for this task and reaches the smaller gap with state-of-the-art models (see SM§11). Notably, adding structured external knowledge or filtered retrieval (as in ViECap or MeaCap) on noise-based decoders improves fluency and informativeness, suggesting such modules are complementary for strong regional semantics. However, they compare similarly to the model using the memory-based decoder.
@@
ZeroCap [59] CVPR’22 - - - - - - - - - - - - - - 14.6 - - - - -
MAGIC [56] ArXiv’22 - - - - - - - - - - - - - - 49.3 - - 17.5 - -
CLOSE [18] ICCV’23 - - - - - - - - - - - - - - 81.2 - - - - -
CapDec [43] EMNLP’22 - - - - - - - - - - - - - - 91.8 - - 35.7 - -
EntroCap [63] NEUCOM’25 - - - - - - - - - - - - - - 94.3 - - 41.5 - -
MERCap [67] AAAI’25 - - - - - - - - - - - - - - 96.0 - - 45.6 - -
ViECap [15] ICCV’23 24.3 74.4 12.0 68.8 14.90 26.4 74.3 15.2 26.6 74.3 102.7 85.0 31.8 74.9 89.7 88.5 75.6 29.8 80.6 70.5
MeaCap [68] CVPR’24 22.5 74.4 12.6 69.8 15.01 28.6 75.1 16.0 28.9 75.1 97.9 85.2 38.6 76.4 86.0 88.6 77.8 40.1 82.8 73.9
DeCap [33] ICLR’23 20.5 75.3 11.2 71.0 17.75 24.6 77.8 17.8 24.9 77.7 95.1 87.4 39.4 78.8 87.4 90.6 79.3 40.0 84.8 76.3
IFCap [31] EMNLP’24 20.8 73.2 10.0 67.4 13.7 21.3 72.5 13.4 22.8 73.0 95.0 83.9 30.5 73.3 86.9 87.3 74.6 29.4 79.2 69.2
Diffusion Bridge [30] CVPR’25 21.8 74.8 10.7 69.8 14.9 20.0 72.8 15.2 20.2 72.8 94.7 85.3 32.9 75.8 85.1 88.8 76.7 33.7 82.1 72.4
@@
RegionCLIP [71] CVPR’22 + Mem. ($\simeq$ DeCap) - - - - 15.85 21.7 76.7 16.01 21.0 75.4 - - - - 93.4 91.2 77.5 38.8 84.5 73.6
AlphaCLIP [57] CVPR’24 + Mem. ($\simeq$ DeCap) 21.3 75.4 11.8 71.0 14.63 19.1 73.9 14.82 19.4 73.8 95.1 87.4 39.5 78.8 89.7 91.1 78.2 40.9 85.4 75.0
@@
Table 2: Comparison of our Patch-ioner framework, using Talk2DINO (T2D), with ZS methods on trace, dense, region-set, and image captioning tasks. Our approach consistently outperforms whole-image and region-supervised baselines in local, fine-grained captioning tasks, while achieving competitive results on whole-image captioning. The table reports CIDEr (C), RefPAC (P), mean average precision (mAP) for dense captioning, and CLIP-Score (CLIP-S) when applicable; best and second-best results are in bold and underlined, respectively.
@@
DeCap: a cat is sleeping on a cluttered desk. Ours (CLIP + Mem.): a cat is sitting at a table with a full laptop. Ours (Talk2DINO + Mem.): office supplies, pens, toys, and other items on desk. GT: Two giraffes, rocks, and a fence. DeCap: a giraffe in a zoo with a city in the background. Ours (CLIP + Mem.): there are some people that are in a lot by a tree. Ours (Talk2DINO + Mem.): two giraffes standing in a fenced area. GT: a clock at a train station. DeCap: a train traveling along the platform of a public train. Ours (CLIP + Mem.): a train is on the tracks and going by. Ours (Talk2DINO + Mem.): a clock on a train station platform above a train. GT: an old man sitting on a bench with a purse. DeCap: a man sitting on a bench while holding a door. Ours (CLIP + Mem.): a bathroom has a blue toilet and the walls. Ours (Talk2DINO + Mem.): a man sits on a wooden bench with a bag on his back. GT: A black cat rubbing against a bottle of wine. DeCap: a black cat standing next to a bottle of wine glasses. Ours (CLIP + Mem.): a cat sitting on the counter of a green bottle. Ours (Talk2DINO + Mem.): a black cat sitting on a chair next to a bottle of wine.
Figure 2: Qualitative results from finer (left) to coarser (right) tasks. Note the discrepancy of predicted and ground-truth captions when an image-level (DeCap, DeCap (Crop)) or a CLIP-based regional (CLIP + Mem.) captioner is applied, with respect to Talk2DINO-based model.